OpenStudy (anonymous):

to solve complete the square, what value should you add to each side of the equation? x^2+10x=-2

Parth (parthkohli):

\[\left(\dfrac{\text{coefficient of }x}{2}\right)^2\]is what you must add to both sides.

OpenStudy (anonymous):

so, it would be 2?

Parth (parthkohli):

no...

OpenStudy (anonymous):

Okay I am not understanding

Parth (parthkohli):

g2g, please ask others for help.

OpenStudy (anonymous):

Ok thanks n e ways

OpenStudy (anonymous):

I got the answer 5^2=25 10/2=25 so the answer 25

OpenStudy (zpupster):

x^2 + 10x + 25 = -2 + 25 (x+5)^2 = 23 take the sqr root both sides (x+5) = + or -sqr root 23 do the math you have 2 answers